A federal regulator is asked to determine the minimum cost reduction required to support the merger of two duopolistic firms. It is important to note that the two firms have similar cost structure with a constant per unit cost and no fixed costs. As of now, each of the two firms produces an output π‘ž = 20 at a price π‘ƒ= 50. 

Note that the socially efficient quantity estimated to be equal to 40. 


You need to find: 

1.     The slope and the vertical intercept of the demand curve

2. The pre-merger average cost. 

3.     The size of the social loss resulting from the merger


1
Expert's answer
2021-03-30T19:55:51-0400

1:slope=change in price/change in quantity

50/20=2.550/20=2.5

Vertical intercept

P=a-b(Q)

P=price

A=intercept where price is 0.

b=slope

Q=quantity

50=aβˆ’2.5(20)50=a-2.5(20)

=100=100

2: 50

3:40βˆ’(20βˆ—2)=040-(20*2)=0